# DFT-s-OFDM-based On-Off Keying for Low-Power Wake-Up Signal

### v1

@@ Abstract
5G-Advanced and likely 6G will support a new low-power wake-up signal (LP-WUS) enabling low-power devices, equipped with a complementary ultra low-power receiver to monitor wireless traffic, to completely switch off their main radio. This orthogonal frequency-division multiplexed (OFDM) signal will emulate an on-off keying (OOK) modulation to enable very low-energy envelope detection at the receiver. Higher rate LP-WUS, containing multiple OOK symbols within single OFDM symbol, will be generated using the time-domain pulse multiplexing of discrete Fourier transform spread (DFT-s-) OFDM. In this context, this paper presents a general signal design framework for DFT-s-OFDM-based OOK generation. General properties of subcarrier coefficients are derived demonstrating that only DFT of the bits needs to be computed online and repeated over the band before applying appropriate frequency-domain processing. The conventional approach of generating rectangular-like OOK waveforms is then addressed by a combination of pre-DFT bit-spreading and post-DFT processing; and the least-squares (LS) method from Mazloum and Edfors, proposed for 5G LP-WUS and also Ambient-IoT, is shown to be implementable as such. Even though aesthetically pleasing and of independent interest, rectangular-like OOK waveforms are not optimal for 5G LP-WUS scenarios due to their limited robustness to channel frequency-selectivity and timing offset, and so shaping methods for spreading the OOK spectrum and concentrating the OOK symbol energy are described and shown to improve the bit error rate performance under practical conditions.
@@ I Introduction
In 4G LTE-M and NB-IoT technologies, 3GPP already defined a WUS to enable receiver sleep modes [1], but its reception relies on a standard OFDM radio, which still demands precise synchronization and high-precision analog-to-digital conversion (ADC). Recently, 3GPP RAN1 completed a study [2] on a so-called low-power WUS (LP-WUS), which is currently under specification for 5G-Advanced. Recent works motivated by this study can be found in [3, 4, 5, 6, 7]. Significantly larger power-saving gains [8] are envisioned if the main radio (MR) of a device could be completely switched off when no message is coming, and being triggered only when necessary. To enable this, the device would need to be equipped with an additional lower-power radio (LR) that monitors possible incoming traffic indicated by an LP-WUS. To achieve ultra low-power consumption, the LR could employ a simple non-coherent envelope detector, through which an LP-WUS can convey information using OOK modulation. Meanwhile, at the transmitter side, an OFDM generation of the LP-WUS is preferred to facilitate implementation and multiplexing with concurrent transmissions. The design of LP-WUS signals is paving the way for other OFDM-based OOK-modulated signals considered in the 5G Ambient-IoT study [9] and likely upcoming 6G IoT technologies.
@@
The problem of generating an OOK-modulated WUS while reusing an existing OFDM implementation has already been considered in the IEEE 802.11ba standard to ease WUS adoption. The IEEE 802.11ba WUS supports a low-rate OOK mode matching the OFDM symbol rate and a high-rate mode, doubling the rate through time-blanking within OFDM symbols. The latter, however, is non-orthogonal to concurrent OFDM signals, leading to inter-carrier interference (ICI). In the context of Wi-Fi, alternative non-orthogonal WUS designs have been proposed in [10, 11], and orthogonal designs in [12, 13, 14]. In [12], an OFDM-based OOK WUS is obtained from numerically-found subcarrier coefficients through minimizing a least-squares (LS) problem with constraints on ON/OFF power levels along with a zero direct current (DC) subcarrier as in legacy Wi-Fi signal format. In the work of Mazloum and Edfors [14], a more general solution is provided by deriving a precoder that maps spread bits to subcarrier coeffients such that it minimizes the LS error from a perfectly rectangular OOK signal. This LS method was one of the main preliminary approaches considered in 3GPP LP-WUS study [15], and is also considered for evaluation purposes in the ongoing 3GPP study on Ambient-IoT [9].
In this paper, we study both the time-domain (TD) and frequency-domain (FD) characteristics of DFT-s-OFDM based OOK modulation, which can enable larger OOK rate than the OFDM symbol rate by exploiting its inherent TD multiplexing property. We are primarily motivated by 5G LP-WUS application in the context of 3GPP standardization, but the provided results are general and may be of independent interest for other applications. The considered system model encompasses an overlaid bit-spreading sequence prior to DFT-precoding followed by spectrum extension (SE) and FD spectral shaping (FDSS). The bit-spreading sequence modulates the DFT-s-OFDM pulses, while SE with FDSS enables control on the number and shape of these pulses [16]. In 3GPP LP-WUS study, defining direct FD subcarrier sequences was considered as an alternative to DFT-s-OFDM with bit-spreading. We link these two methods by showing that a common overlaid sequence on ON-symbols can equivalently be implemented as an FD overlaid sequence. This provides an equivalent low-complexity implementation where the DFT precoder size needs only to match the number of bits – i.e., without bit spreading. Moreover, compact DFT formulas and tables absorbing Manchester encoding are obtained for efficient storage.
@@
Next, we focus on creating rectangular-like OOK waveforms, as considered in IEEE 802.11ba WUS and the early stages of the 3GPP LP-WUS study. We employ a linear phase ramp as the bit-spreading sequence that minimize the coherent combining of consecutive DFT-s-OFDM pulses, thereby flattening the envelope of ON symbols. We also discuss rectangular-like OOK waveforms with the additional feature of a zero DC component. Furthermore, we show that the waveform resulting from ‘LS precoding’ of [14], treated in [15] as an alternative to DFT precoding, can be obtained as a specific implementation of DFT-s-OFDM-based OOK.
@@
Rectangular-like OOK waveforms, which may be relevant to small indoor coverage scenarios like in WiFi and Ambient-IoT, are not suitable for 5G LP-WUS. Large cell coverage, as supported in 3GPP standards, involves broader bandwidth usage, making it prone to greater frequency-selectivity and channel time dispersion. Additionnally, 3GPP considered a significantly higher sampling rate at LR compared to the transmitted OOK symbol rate. In this context, as long as the OOK signal achieves a sufficient ON/OFF energy split, envelope fluctuations actually matter little. Relaxing the rectangular waveform target, shaped OOK providing improved detection performance are discussed. We formalize the effect of phase scrambling in the overlaid sequence on the power spectrum distribution, as considered in [15] to flatten the spectrum. Additionally, robustness to timing errors is achieved by concentrating energy within the ON symbols. Finally, simulation results are provided to confirm the benefits of such shaping over rectangular OOK.
@@
The remainder of the paper is as follows. Section II introduces the LP-WUS system model from an OFDM transmitter to an envelope detector at the LR. Section III describes the OOK signal generation framework based on DFT-s-OFDM modulation highlighting its general TD and FD properties. Section IV addresses the generation of rectangular-like OOK waveform, while Section V discussed shaping methods relevant for mitigating wireless channel impairments. Numerical evaluations are presented in Section VI, and conclusions are drawn in Section VII.
@@ II-A OFDM Modulation
The LP-WUS is generated by populating $N_{\rm sc}$ subcarriers within other data symbols, all multiplexed by an OFDM modulation using an $N_{\rm fft}$-point IFFT followed by cyclic prefix (CP) addition. While such setup avoid interference from the OOK WUS to concurrent data received by OFDM terminals, the WUS itself is intended for a non-OFDM, envelope detector based, receiver for which concurrent data may interfere and thus some guard bands may still be necessary. The total LP-WUS bandwidth allocation is thus $(N_{\rm sc}+2N_{\rm GB})$ where $N_{\rm GB}$ are null subcarriers on each band side.
@@ II-B Envelope Detection
The receiver is considered to capture the WUS and noise power over the WUS bandwidth, including guard bands, such that the SNR is defined as $\mathsf{snr}=\frac{E_{h}P_{W}}{(N_{\rm sc}+N_{\rm GB})N_{0}}$ where $E_{h}=\sum_{p=0}^{L_{p}-1}E_{p}$ is total average channel energy, and $P_{W}=\operatorname{E}\left[|s^{W}[n]|^{2}\right]=\sum_{k=0}^{N_{\rm sc}-1}\operatorname{E}\left[|X^{W}[k]|^{2}\right]$ is the average transmitted power of the WUS signal.
@@ III Generation Framework for DFT-s-OFDM-based OOK
A generic method to construct LP-WUS subcarrier coefficients $\{X^{W}[k]\}_{k=0}^{N_{\rm sc}-1}$ is to use DFT precoding, resulting in a DFT-s-OFDM modulation only for $s^{W}[n]$ but not for the remaining part of the OFDM signal $s^{D}[n]$ This is the retained approach by 3GPP [2].
@@ III-A DFT-s-OFDM-based OOK
The considered DFT-s-OFDM based OOK generation framework for LP-WUS is schematized in Figure 2.
@@ III-A1 Input Bits and Manchester Coding
First, a bit string $b[l]\in\{0,1\}$, $l=0,\ldots,N_{\rm bit}-1$ is taken as input. We will consider that this bit string follows from a Manchester encoding of an original bit string $b_{o}[n]$, $n=0,\ldots,N_{\rm bo}-1$ such that $N_{\rm bo}=N_{\rm bit}/2$ and
@@ III-A2 Bit Spreading
The bits $b[l]$ are spread by a factor $N_{\rm seg}=N_{\rm pulse}/N_{\rm bit}$ to obtain a sequence of symbols $d[m]$ that will be used to modulate $N_{\rm pulse}\leq N_{\rm sc}$ DFT-s-OFDM pulses. If a bit $b[l]=0$ then it is mapped to an all-zero sequence; otherwise if $b[l]=1$ it is mapped to a (so-called in 3GPP) overlaid sequence $\tilde{r}_{l}[n]$, $n=0,\ldots,N_{\rm seg}-1$. In general, each bit $b[l]$ can be written as $N_{\rm seg}$-time repeated as $b_{\rm r}[m]=b[l_{m}]$ with $l_{m}=\lfloor\frac{m}{N_{\rm seg}}\rfloor$ and element-wise multiplied with a concatenated $N_{\rm pulse}$-long bit-spreading sequence $r[m]=\tilde{r}_{l_{m}}[m-N_{\rm seg}l_{m}]$. The modulation symbols are thus

### v2

@@ Abstract
5G-Advanced and likely 6G will support a new low-power wake-up signal (LP-WUS) enabling low-power devices, equipped with a complementary ultra low-power receiver to monitor wireless traffic, to completely switch off their main radio. This orthogonal frequency-division multiplexed (OFDM) signal will emulate an on-off keying (OOK) modulation to enable very low-energy envelope detection at the receiver. Higher rate LP-WUS, containing multiple OOK symbols within single OFDM symbol, will be generated using the time-domain pulse multiplexing of discrete Fourier transform spread (DFT-s-) OFDM. In this context, this paper presents a comprehensive signal design framework for DFT-s-OFDM-based OOK generation. General properties of subcarrier coefficients are derived demonstrating that only DFT of the bits needs to be computed online and repeated over the band before applying appropriate frequency-domain processing. The conventional approach of generating rectangular-like OOK waveforms is then addressed by a combination of pre-DFT bit-spreading and post-DFT processing; and the least-squares (LS) method from Mazloum and Edfors, proposed for 5G LP-WUS and also Ambient-IoT, is shown to be implementable as such. Even though aesthetically pleasing and of independent interest, rectangular-like OOK waveforms are not optimal for 5G LP-WUS scenarios due to their limited robustness to channel frequency-selectivity and timing offset, and so shaping methods for spreading the OOK spectrum and concentrating the OOK symbol energy are analyzed and shown to improve the bit error rate performance under practical conditions.
@@ I Introduction
In 4G LTE-M and NB-IoT technologies, 3GPP already defined a WUS to enable receiver sleep modes [1], but its reception relies on a standard OFDM radio, which still demands precise synchronization and high-precision analog-to-digital conversion (ADC). Recently, 3GPP RAN1 completed a study [2] on a so-called low-power WUS (LP-WUS) with a first specification for 5G-Advanced. Recent works motivated by this study can be found in [3, 4, 5, 6, 7]. Significantly larger power-saving gains [8] are envisioned if the main radio (MR) of a device could be completely switched off when no message is coming, and being triggered only when necessary. To enable this, the device would need to be equipped with an additional lower-power radio (LR) that monitors possible incoming traffic indicated by an LP-WUS. To achieve ultra low-power consumption, the LR could employ a simple non-coherent envelope detector, through which an LP-WUS can convey information using OOK modulation. Meanwhile, at the transmitter side, an OFDM generation of the LP-WUS is preferred to facilitate implementation and orthogonal multiplexing with concurrent transmissions. The design of LP-WUS paved the way for other OFDM-based OOK-modulated signals in 3GPP, already for the OOK signal in Ambient-IoT (A-IoT) [9] and likely for other upcoming 6G IoT technologies.
@@
Background on OOK for Wi-Fi WUS: Generating an OOK-modulated WUS while reusing an existing OFDM implementation has already been considered in the IEEE 802.11ba Wi-Fi standard in order to ease WUS adoption. The Wi-Fi WUS supports a low-rate OOK mode matching the OFDM symbol rate and a high-rate mode, doubling the rate through time-blanking within OFDM symbols. The latter, however, is non-orthogonal to concurrent OFDM signals, resulting in inter-carrier interference (ICI). In the context of Wi-Fi, non-orthogonal WUS designs have been proposed in [10, 11], and orthogonal OFDM-based designs in [12, 13, 14]. In [12], sequences of subcarrier coefficients are numerically-found under several constraints such as ON/OFF power levels, and a zero direct current (DC) subcarrier as in legacy Wi-Fi signal format. In [12], Golay sequences are used instead to limit the high Peak-to-Average Power Ratio (PAPR) from OFDM multiplexing of multiple, but low-rate, OOK signals. A scalable high-rate OFDM-based approach is proposed by Mazloum and Edfors in [14] by deriving a precoder, mapping spread bits to subcarrier coefficients, such that the resulting OFDM signal minimizes the least-square (LS) error from a perfectly-rectangular OOK signal.
Background on OOK for 5G-Advanced LP-WUS: When preparing our contributions to 3GPP standardization of LP-WUS, our initial approach independently explored the inherent time-domain (TD) pulse multiplexing property of DFT-s-OFDM to generate higher rate but OFDM-compliant OOK signals. We first targeted a rectangular-like OOK by controlling pulse combining along with frequency-domain (FD) spectral shaping (FDSS) as well as spectrum extension (SE), and later on investigating other shaping techniques for performance improvement, including FD repetition, alternative bit-spreading sequences and TD pulse shaping. This is reflected in related 3GPP contributions, e.g., [15, 16, 17] which discuss the LS method from [14], ±$\pm 1$ alternating bit-spreading sequence for rectangular OOK, Zadoff-Chu (ZC) bit-spreading for improving performance in frequency-selective channels, and concentrated OOK for robustness against timing misalignment. These signal designs were also captured in [2] as potentially useful for performance improvement.
From the perspective of wireless standard engineers, using DFT-s-OFDM to create OFDM-compliant OOK signals is a rather natural approach, and similar considerations were therefore independently proposed by many 3GPP participants from the beginning of the 3GPP LP-WUS study [18]. In retrospect, we identify the possibly first occurrence of DFT-s-OFDM-based OOK in a Wi-Fi WUS standardization contribution by A. Sahin, R. Yang et al. [19], where, interestingly, FDSS-SE and ZC bit-spreading are already listed among possible signal design ingredients. However, as is often the case, standardization contributions focus on addressing specific use cases with a limited scope and less detailed analysis of signal design alternatives – for instance the choice of a ZC sequence in [19] is not further justified.
The LS processing from [14] to obtain rectangular-like OOK, introduced in 3GPP by us [15] and others [18], was also one of the main approaches considered at the initial stage of the LP-WUS study. As reflected in [2], this approach was seen by many participants as a specific OOK generation method, involving a specific and so-called “LS precoding”, distinct from the DFT precoding of DFT-s-OFDM – even though this LS precoder is simply a larger, truncated DFT matching both the OFDM sampling rate at its input and the bandwidth at its output. Direct FD sequence-based OFDM-OOK was also suggested in 3GPP LP-WUS study as an alternative to TD approaches. Eventually, 5G LP-WUS is a DFT-s-OFDM-based OOK signal with ZC overlaid sequence.
Background on OOK for 5G-Advanced A-IoT: LP-WUS study influenced the more recent 3GPP study on A-IoT [9] for the reader-to-device OOK waveform generation. In this context, a rectangular-like OOK signal design can be a relevant approach notably as edge/transition detection between adjacent OOK symbols is likely to be used at receiving devices, for which OOK symbol flatness becomes a desired signal property [20, 21]. Both DFT-s-OFDM-based and LS approaches have been considered for A-IoT [22] and enabled by a signal generation that allows a DFT precoder larger than the subcarrier allocation, unlike LP-WUS. Other waveform generation steps are left to manufacturer-specific implementation choices.
Motivations, Benefits and Limitations: In both LP-WUS and A-IoT, the primary motivation of using DFT-s-OFDM for OOK lies in its time-multiplexing property, while still complying with legacy OFDM-based system and hardware, and not for its traditionally considered main benefit which is a low PAPR. Notably, these applications consider OOK transmission inserted in a downlink OFDM signal whose other subcarriers could carry high-order QAM symbols, resulting overall in a high PAPR. In this context, the transmitted signal as a whole is OFDM-based and not DFT-s-OFDM-based.
When DFT-s-OFDM is used in 5G uplink data transmission, and with $\pi/2$-BPSK, it can be complemented by FDSS to achieve very low PAPR. For higher order QAM, FDSS with SE has been also considered [23]. The low-PAPR benefit of FDSS-SE in this context comes at the cost of breaking orthogonality among the DFT-s-OFDM multicarrier pulses, thus creating ICI among the data modulation symbols. For OOK signal design, FDSS-SE can similarly enable control over the signal’s envelope, but in this case, without any interference issue as OOK detection is blind to the orthogonality property of the underlying DFT-s-OFDM pulses. On the contrary, by limiting energy leakage among consecutive OOK symbols, FDSS-SE can help to mitigate interference among them. In any case, the OFDM subcarriers of the OOK signal, with or without FDSS-SE, remain orthogonal to those of concurrent transmissions and thus interference-free to them.
Another concern that could arise from OOK – especially rectangular OOK – is the potentially large out-of-band emission (OOBE). However, with an OFDM-based approach, most of the power is by design concentrated within the subcarrier allocation, yielding an OOBE level comparable to typical OFDM signals [19]. The OOBE of OFDM is well known to not directly satisfy regulatory requirements, but standard filtering techniques such as root-raised-cosine filter can often address this issue effectively. Moreover, in practice this OOK signal may be inserted within an OFDM signal of larger bandwidth, rendering its impact negligible on the overall OOBE. On the other hand, an inherent limitation of the OFDM-based approach is that the transition slopes between OOK symbols are constrained by the subcarrier allocation.
This Paper’s Contributions: We provide a comprehensive framework for DFT-s-OFDM-based OOK modulation and study both its TD and FD characteristics. While primarily motivated by the 5G LP-WUS application, the presented results are general and may be of independent interest for other use cases. The considered framework encompasses an overlaid bit-spreading sequence prior to DFT-precoding, possibly combined with FDSS and SE. The bit-spreading sequence modulates the DFT-s-OFDM pulses, while FDSS-SE offers additional degrees of freedom for shaping the pulses and controlling their number.
In both Wi-Fi WUS and the 5G LP-WUS study, defining subcarrier sequences directly was also considered as an alternative to DFT-s-OFDM with bit spreading. We link these two methods by showing that a common TD overlaid sequence on top of bit spreading results in an FD overlaid sequence that creates an ON-symbol kernel whose TD multiplexing is controlled by the DFT of the bits. This insight reveals an equivalent low-complexity implementation, where the data-dependent DFT precoding is applied directly to the input bit string – without the need for explicit bit spreading. Furthermore, compact DFT formulas and tables that absorb Manchester encoding are derived for efficient storage.
@@
Next, we focus on the design of rectangular-like OOK waveforms, as considered in Wi-Fi WUS, the early stage of the 5G LP-WUS study, and 5G A-IoT study. A linear phase ramp is applied in the bit-spreading sequence to minimize the coherent combining of consecutive DFT-s-OFDM pulses, thereby flattening the ON-symbol envelope. We also demonstrate how the bit-spreading sequence can be chosen to obtain a rectangular-like OOK waveform with the additional feature of a zero DC component. Moreover, we show that the LS waveform from [14] arises as a special case of the considered DFT-s-OFDM-based OOK modulation, using a specific FDSS window.
@@
Finally, we move beyond the intuitive approach of regular and rectangular-like OOK waveform to propose and analyze shaping methods better suited for LP-WUS application scenarios. Rectangular-like OOK designs, though relevant to small indoor coverage applications such as Wi-Fi and A-IoT, are not optimal for 5G LP-WUS. Large cell coverage, as supported by the 5G NR standard, involves larger bandwidth usage and increased sensibility to frequency-selectivity and channel time dispersion. Additionally, 3GPP considered a significantly higher sampling rate at LR compared to the transmitted OOK symbol rate. In this context, as long as the OOK signal achieves a sufficient ON/OFF energy split, envelope fluctuations actually matter little.
We formalize the effect on the power spectrum distribution obtained from phase scrambling by an ZC overlaid sequence, as considered in [24] to flatten the spectrum. More precisely, we show that this results in a constant spectral comb over the bandwidth, containing half the total signal power and thereby providing excellent frequency diversity. Additionally, we investigate the TD concentration of ON-symbol energy to enhance robustness against timing errors. Simulation results confirm the performance benefits of such shaping methods over rectangular OOK in LP-WUS application.
@@
Organization: The remainder of the paper is as follows. Section II introduces the LP-WUS system model from an OFDM transmitter to an envelope detector at the LR. Section III describes the OOK signal generation framework based on DFT-s-OFDM modulation highlighting its general TD and FD properties. Section IV addresses the generation of rectangular-like OOK waveform, while Section V discussed shaping methods relevant for mitigating wireless channel impairments. Performance evaluations are presented in Section VI, and conclusions are drawn in Section VII.
@@ II-A OFDM Modulation
The LP-WUS is generated by populating $N_{\rm sc}$ subcarriers within other data symbols, all multiplexed by OFDM using an $N_{\rm fft}$-point IFFT followed by cyclic prefix (CP) addition. While such setup avoids interference to concurrent data received by OFDM terminals, the WUS itself is intended for a non-OFDM, envelope-detection receiver for which concurrent data may interfere, and thus some guard bands may still be necessary. The total LP-WUS bandwidth allocation is thus $(N_{\rm sc}+2N_{\rm GB})$ where $N_{\rm GB}$ are null subcarriers on each band side.
@@
A CP protects OFDM symbols from interfering on each other after channel dispersion. From the OOK signal perspective only, CP insertion is mainly detrimental, with its presence and length imposed on LP-WUS for system compliance. The CP reduces the OOK transmission rate while offering no interference protection when more than one OOK symbol is transmitted per OFDM symbol. To avoid this rate reduction, the targeted OOK signal is punctured in [14], thereby maintaining a regular bit period even after CP insertion, rendering LP-WUS synchronization independent of the OFDM structure, but at the cost of increased errors in bit detection. Moreover, CPs can introduce false rising/falling edges – a concern raised in the A-IoT study [9]. In both LP-WUS and A-IoT, the main CP handling approach assumes that the receiving device removes it without requiring specific transmitter-side assistance. For high-rate OOK, where CP becomes comparable to or longer than an OOK symbol, receiver-only handling may become insufficient, and transmitter-side handling – such as reserving the last OOK symbols to control the CP, as discussed in [9] – may also be needed.
Figure 1: Illustration of DFT-s-OFDM based OOK modulation
@@ II-B Envelope Detection
The receiver is considered to capture the WUS and noise power over the WUS bandwidth, including guard bands, such that the SNR is defined as $\mathsf{snr}=\frac{E_{h}P_{W}}{(N_{\rm sc}+2N_{\rm GB})N_{0}}$ where $E_{h}=\sum_{p=0}^{L_{p}-1}E_{p}$ is total average channel energy, and $P_{W}=\operatorname{E}\left[|s^{W}[n]|^{2}\right]=\sum_{k=0}^{N_{\rm sc}-1}\operatorname{E}\left[|X^{W}[k]|^{2}\right]$ is the average transmitted power of the WUS signal.
@@ III Generation Framework for DFT-s-OFDM-based OOK
In each OFDM symbol, the LP-WUS OOK signal contains $N_{\rm bit}$ OOK symbols (ON or OFF), identified as $N_{\rm bit}$ coded bits being the line coding of $N_{\rm bo}$ info bits. A generic method to achieve a multi-bit OFDM-OOK for LP-WUS, i.e. $N_{\rm bit}>1$ bits per OFDM symbol, is to obtain the subcarrier coefficients by DFT of a string of $N_{\rm pulse}\leq N_{\rm sc}$ modulation symbols that embeds the targeted ON/OFF pattern, resulting in a DFT-s-OFDM modulation only for $s^{W}[n]$, but not for the remaining part of the OFDM signal $s^{D}[n]$. As a result, the $N_{\rm pulse}$ input symbols then modulates or deactivates a set of selected DFT-s-OFDM time pulses within the OFDM symbol.
@@ III-A DFT-s-OFDM-based OOK
The considered DFT-s-OFDM based OOK generation framework for LP-WUS is schematized in Fig. 1.
@@ III-A1 Input Bits and Manchester Coding
The ON/OFF pattern in the OFDM symbol is controlled by the coded bit string $b[l]\in\{0,1\}$, $l=0,\ldots,N_{\rm bit}-1,$ which we will consider (most of the time) to be the Manchester encoding of an original info bits string $b_{o}[n]\in\{0,1\}$, $n=0,\ldots,N_{\rm bo}-1$ of length $N_{\rm bo}=N_{\rm bit}/2$. This Manchester encoding can be formally written as
@@ III-A2 Bit Spreading
The bits $b[l]$ are spread by an integer factor $N_{\rm seg}=N_{\rm pulse}/N_{\rm bit}$ to obtain a sequence of symbols $d[m]$ that will be used to modulate $N_{\rm pulse}\leq N_{\rm sc}$ DFT-s-OFDM pulses. If a bit $b[l]=0$ then it is mapped to an all-zero sequence; otherwise if $b[l]=1$ it is mapped to a (so-called in 3GPP) overlaid sequence $\tilde{r}_{l}[n]$, $n=0,\ldots,N_{\rm seg}-1$. In general, each bit $b[l]$ can be written as $N_{\rm seg}$-time repeated as $b_{\rm r}[m]=b[l_{m}]$ with $l_{m}=\lfloor\frac{m}{N_{\rm seg}}\rfloor$ and element-wise multiplied with a concatenated $N_{\rm pulse}$-long bit-spreading sequence $r[m]=\tilde{r}_{l_{m}}[m\;{\rm mod}\;N_{\rm seg}]$. The modulation symbols are thus